pls help to to answer: what is the lenght of the wire where it cut into the ratio of 2:3:5 when its midle part is 7 m?
Answer by ptaylor(2198)

Let L=length of the wire
Now we know that for a 10ft length of wire, for example, the three pieces will be 2, 3 and 5 feet, respectively, soooo:
(2/10)L=first part
(3/10)L=second part
(5/10)L=third part
We are told that the middle part is 7 m. Our equation to solve, then, is:
(3/10)L=7
L=70/3=23 1/3 m
In other words if the ratio of cats, to dogs, to horses is 2:6:7 this simply means that out of every 2+6+7=15 animals, 2/15 will be cats, 6/15 will be dogs and 7/15 will be horses.
Problem: If the ratio of girls to boys in a class is 5:6, How many girls are in a class of 66?
Answer: (5/(6+5))*66=(5/11)*66=30 girls
Hope this helps---ptaylor